OpenStudy (anonymous):

Find the equation of a parabola with the vertex at (-3,-2) and directrix of X = 1

OpenStudy (loser66):

(y+2)^2= 4(x+3)^2

OpenStudy (loser66):

I think so

OpenStudy (anonymous):



OpenStudy (anonymous):

y = (x+3)^2 - 2

OpenStudy (loser66):

@Euler271 can you explain me how to get that equation?

OpenStudy (anonymous):

you were thinking the equation of a circle. y = a(x-h)^2 + k is the general equation of a parabola, where the point (h, k) is the vertex and a is the directrix.

OpenStudy (loser66):

nope, y ^2 = 4px with directrix x = 1( or p = 1) is it not from conic section?